An electron is released from rest in a uniform electric field and accelerates to the north at a rate of 265 m/s2 .

What is the magnitude of the electric field?

Express your answer to three significant figures and include the appropriate units.
Answer #1

as we know

F = qE = ma

1.6*10^-19* E = 9.11*10^-31* 265

E = 1.509*10^-9 N/C

=======

Comment in case any doubt, will reply for sure . Goodluck